### v3

# Quantum oscillation study of the large magnetoresistance in Mo substituted WTe2 single crystals

Abstract
The list of interesting electrical properties exhibited by transition metal dichalcogenides has grown with the discovery of extremely large magnetoresistance (MR) and type-II Weyl semimetal behaviour in WTe2 and MoTe2. The extremely large MR in WTe2 is still not adequately understood. Here, we systematically study the effect of Mo substitution on the quantum oscillations in the MR in WTe2. The MR decreases with Mo substitution, however, the carrier concentrations extracted from the quantum oscillations show that the charge compensation improves. We believe that earlier interpretations based on the two-band theory, which attribute the decrease in MR to charge imbalance, could be incorrect due to over-parametrization. We attribute the decrease in MR in the presence of charge compensation to a fall in transport mobility, which is evident from the residual resistivity ratio data. The quantum scattering time and the effective masses do not change within experimental errors upon substitution.
pacs:
I Introduction
With the rise of graphene, other two-dimensional materials have emerged, among which the transition metal dichalcogenides (TMDCs) have proved to be particularly notable [1, 2]. TMDCs such as MoS2 offer enormous potential in photonics and optoelectronic devices, while others, including NbSe2 and TaSe2, display interesting many-particle electronic states like charge density waves and superconductivity [3]. TMDCs have received renewed interest after various unique properties were recently discovered in WTe2 and MoTe2. These include the discovery in WTe2 of an extremely large non-saturating magnetoresistance (MR) [4] and superconductivity at high pressures [5, 6], and in MoTe2 of a large MR [7, 8] and an enhancement of the superconducting transition temperature under pressure [9]. In addition, WTe2 is predicted to be a type-II Weyl semimetal [10] and this aspect has been probed successfully by angle resolved photoelectron spectroscopy (ARPES) [11, 12, 13]. Similar studies have also confirmed the type-II Weyl character of MoTe2 [14, 15] and Mo substituted WTe2 (MoxW1-xTe2) [16, 17, 18]. Monolayer WTe2 is also reported to be a quantum spin Hall insulator [19, 20] and charge-to-spin conversion [21] as well as edge transport have been reported [22]. Quantum oscillations and high-mobility electrons have recently been reported in topologically insulating monolayers of WTe2 [23].
The extremely large MR in WTe2 has been attributed to nearly perfect compensation of electrons and holes in the two-band model [4]. ARPES and quantum oscillations, which can give information about the carrier concentrations, mostly report four pockets; two electron and two hole pockets [24, 25, 26, 27, 28, 29, 30]. Proper mapping of the frequencies in the quantum oscillations with electron or hole pockets is vital in determining whether the electron and hole concentrations are equal. Most studies attribute the smallest and largest frequencies to hole pockets and the middle two frequencies to electron pockets. However, Cai et al. [31] have identified them differently in a pressure-dependent study.
While the MR is typically discussed in terms of charge compensation, alternative mechanisms, like a field-induced metal-insulator transition [32, 29] and Kohler’s law [33], have also been proposed. Moreover, in the two-band model, there is no consensus on whether the mobility or the charge compensation is more important in accounting for the observed behaviour [34, 35, 36, 37, 30]. Often, the MR remains large and non-saturating even when a significant charge imbalance is introduced via gating [38]. The MR decreases at higher temperatures, which is attributed to a deviation from charge compensation. However, an ARPES study has found no evidence of such a deviation in the electronic structure [39] while another ARPES study found that the charge compensation is possible at best only in a very narrow range of temperature [40]. All this has raised questions about the appropriateness of the two-band model for the MR in WTe2. Studies on the effect of non-stoichiometry caused by Te vacancies and Mo substitution (MoxW1-xTe2) can shed further light on the large MR in WTe2. Although these studies report a decrease in MR with Mo substitution, the carrier concentrations extracted from the quantum oscillations have not previously been compared to those extracted from the fit to the two-band theory [41, 42, 43]. Here, we study the quantum oscillations and MR of four different compositions to shed light on the evolution of the Fermi surfaces and carrier concentrations with Mo substitution in WTe2, and test the validity of the charge compensation theory.
II Experimental Details
Single crystals of various compositions of Mo substituted WTe2 (W1-xMoxTe2, , 0.05, 0.1, 0.2) were grown using chemical vapour transport technique with TeCl4 as the transport agent. For pure WTe2, the starting material was W and Te powder in stoichiometric ratio, while for the substituted compounds it was polycrystalline powder prepared earlier by heating stoichiometric ratio of the constituent elements W, Mo and Te at 700-760 C for 5-7 days in evacuated quartz tubes. For the crystal growth, the starting material was taken in 20-25 cm long quartz tubes and evacuated and sealed. The tubes were then placed in horizontal three zone furnaces with the charge end and the growth end of the tubes kept at 900 C and 700-800 C, respectively. The growth was performed over a period of three weeks. X-ray diffraction measurements were performed on the single crystals using a high resolution Panalytical X’Pert Pro diffractometer with Cu K radiation ( Å) and Cu K hybrid monochromator. The crystals were mounted on zero diffraction plates made of silicon cut at special orientation so that there were no background Bragg peaks from the sample holder. A Laue X-ray imaging system with a Photonic-Science Laue camera was used to acquire the Laue back-scattered diffraction images of the single crystals. The chemical composition of the single crystals was determined using an Oxford Instruments energy dispersive X-ray spectrometer fitted in a Zeiss SUPRA 55–VP scanning electron microscope. Electrical transport and MR measurements down to 2 K and in magnetic fields up to 9 T were performed in a Quantum Design Physical Property Measurement System (PPMS) using the AC transport measurement option. Four probe electrical contacts were made on the single crystals using silver wires and conductive silver paint. Currents of 1 mA were used in the AC resistivity measurements.
III Results
Fig. 1(a) shows the representative image of the single crystals of W1-xMoxTe2 (), grown by the chemical vapour transport (CVT) method. WTe2 grows in the distorted octahedral configuration, known as T-WTe2 [44, 45], with the space group [46]. Fig. 1(b) shows the (008) Bragg peak seen in the X-ray diffraction (XRD) pattern for the four compositions. The lattice parameter , obtained from the XRD data, decreases with increasing Mo substitution, as seen in Fig. 1(c). Since the radius of Mo4+ ion (0.65 Å) is smaller than that of the W4+ ion (0.66 Å) [41], the lattice parameter along the -axis decreases with Mo substitution and the Bragg peaks shift towards higher values. The value for obtained for the non-substituted WTe2 single crystal is Å which is in agreement with that reported in literature [46, 47].
Figure 1(d) shows the Laue back-scattered diffraction image along the [001] direction for a crystal with the composition W0.95Mo0.05Te2 and demonstrates the very high quality of the crystals. The compositions of the substituted W1-xMoxTe2 crystals were determined using energy dispersive X-ray spectroscopy (EDS) and were found to be close to their nominal values.
| Composition | RRR | MR |
| WTe2 | 79 | 8885 |
| W0.95Mo0.05Te2 | 47 | 4871 |
| W0.9Mo0.1Te2 | 30 | 1996 |
| W0.8Mo0.2Te2 | 24 | 700 |
Electrical resistivity and MR measurements were performed on single crystal samples of the four compositions of W1-xMoxTe2. Fig. 2 shows the temperature variation of the normalised resistivity from room temperature down to 2 K. It is clear that all the compositions studied preserve the metallic nature of WTe2, implying that the T structure of WTe2 is retained for these levels of Mo substitution. This is in agreement with the published phase diagram of the WTe2-MoTe2 system [48, 49]. The residual resistivity ratio (RRR) defined as for each of the four compositions is summarized in Table 1. The RRR value for our pure WTe2 sample is comparable to that previously reported for WTe2 single crystals grown by CVT [28]. The RRR decreases with increasing Mo concentration in keeping with previous studies [41]. This is a sign of increased scattering at low temperatures due to an increase in the density of defects with substitution leading to a decrease in the transport scattering time and hence also the mobility.
The MR of the single crystal samples of each of the four compositions was measured at low temperatures and in magnetic fields up to 9 T. All MR measurements were performed with the magnetic field applied parallel to the -axis, easily identified as the direction perpendicular to the flat surfaces of the crystals. Fig. 3 shows the percentage relative change in MR at a few representative temperatures of 2, 3 and 5 K for the four compositions.
The relative change in MR is defined as,
where is the longitudinal resistivity and is the magnetic field. The values of % relative change in MR at 9 T are given in Table 1. A large non-saturating MR is seen in all four different compositions, with a maximum of nearly 9000% in case of the unsubstituted WTe2 sample. The relative change in MR shows a trend of decreasing magnitude as one goes from the unsubstituted WTe2 to the 20% Mo substituted WTe2 sample, which we propose is caused by the decreasing transport mobility. The MR also decreases as temperature increases for all the compositions.
All four compositions show prominent Shubnikov-de Haas oscillations in the MR. The oscillatory component of the MR is isolated by subtracting a second-order polynomial background from the raw data. Fig. 4 shows these extracted quantum oscillations for the four different compositions at 2 K. The individual oscillation frequencies correspond to extremal cross-sectional areas of different Fermi surfaces (electron or hole pockets) and are obtained by performing a fast Fourier transform (FFT) of an average of the oscillations in both rising and falling fields for T.
| Carrier concentration | WTe2 | W0.95Mo0.05Te2 | W0.9Mo0.1Te2 | W0.8Mo0.2Te2 | |
| () | This work | Previous work [28] | |||
| 5.7 | |||||
| 8.4 | |||||
| 10 | |||||
| 11.14 | - | ||||
The results of the FFT for the pure sample at 2 K [see Fig. 5(a)], show four peaks at the frequencies T, T, T and T. These agree very well with those previously reported in literature for WTe2 [28, 27, 31]. We attribute the smallest and largest frequencies ( and ) to a pair of nearly concentric hole pockets and the two middle frequencies ( and ) to a pair of concentric electron pockets, following Refs. 28, 27, which rely on band structure calculations to match the frequencies with hole and electron pockets.
The frequencies of the quantum oscillations at 2 K for the four compositions as a function of Mo substitution are plotted in Fig. 6. These values are also given in Supplemental Material Table S1 [50] along with previously reported values for unsubstituted WTe2. While and increase somewhat with increased Mo content, and remain largely stable. The peak for the frequency is not resolved in the FFT of the oscillations in W0.8Mo0.2Te2.
The carrier concentration of the hole or electron pockets can be obtained via the Onsager relation [27, 51]
which relates the frequency of oscillation F to the extremal cross-sectional area of the Fermi surface in a direction perpendicular to the applied field. Previous studies suggest that the Fermi pockets in WTe2 are triaxial ellipsoids [27, 28]. However, angle-dependent quantum oscillation [27] and anisotropic MR studies [52] have shown only mild anisotropy of the Fermi surface. The ratio of the Fermi wavevector () along the and directions () is expected to be less than 2.5 for all the four pockets [27]. This is much less than seen in other 2D materials, e.g. graphite for which [52]. The mild anisotropy has been used to justify the simplifying assumption of a spherical Fermi surface i.e. [29, 41] for WTe2 which then allows an estimate to be made of the carrier density .
Td-MoTe2 and Td-Mo1-xWxTe2 are isostructural to WTe2 and their the lattice parameters do not differ much [53, 54], suggesting that Mo substitution affects the lattice parameters isotropically, from which it can be assumed that the Fermi surface would also be affected isotropically. Indeed, ARPES studies have confirmed that the Fermi-surface topology in Mo1-xWxTe2 is unchanged from that of WTe2 [53, 54], implying that if a spherical Fermi surface can be assumed for WTe2 then it can also be assumed for Mo substituted WTe2. We have calculated the carrier concentrations for each of the four frequencies for the compositions using the spherical surface assumption and these are presented in Table 2.
Two of each Fermi-surface pocket is known to be present within the Brillouin zone, one on each side of the point [27, 28, 29]. As a result, the total electron concentration and the total hole concentration and their values are given in Table 3. The electron concentration is slightly higher than the hole concentration in the unsubstituted sample. However, with increasing Mo substitution, the oscillation frequencies change as described above and hence the hole concentration increases slightly while the electron concentration remains the same. This would imply that the charge compensation actually becomes better with Mo substitution. The slightly different effect of the Mo substitution on the electron and hole pockets is likely due to different curvatures of the bands.
Figure 7 shows the temperature dependence of the FFT amplitudes of each of the four peaks observed in the FFT spectra. The solid lines are fits to the Lifshitz-Kosevich formula , where with T being the the average of the inverse magnetic-field window of the FFT [55, 51]. From this we extract the effective masses given in Supplemental Material Table S2 [50] and plotted in Fig. 8. There is not much change in the effective masses with Mo doping, further indicating that Mo doping does not affect the Fermi surface topology.
By performing a Dingle analysis, the quantum oscillations can also be used to find the quantum scattering time (), which is different from the transport scattering time () [56]. It is the time an electron spends in a particular Bloch state and is affected equally by both small and large-angle scattering events. On the other hand, the transport scattering time represents the time between two collisions in electron transport. It is weighted by a factor where is the angle of scattering. Thus, it is affected more by large-angle scattering than by small-angle scattering.
We perform the Dingle analysis [51] for the pocket corresponding to the frequency for all four compositions and the relevant plots are shown in Fig. 9. From the linear fit to the data, the Dingle temperature, and the quantum scattering time , which are related according to are obtained for the holes corresponding to the Fermi-surface pocket with frequency and the values are given in Table 3. Interestingly, the quantum scattering times obtained from the Dingle analysis do not show significant change across the four compositions, whereas a decrease in transport mobility implies a decrease in the transport scattering time. Significant differences between transport and quantum scattering times are expected and are seen, for example, in two-dimensional electron gases and graphene [57, 58].
| Composition | WTe2 | W0.95Mo0.05Te2 | W0.9Mo0.1Te2 | W0.8Mo0.2Te2 | |
| This work | Previous work | ||||
| (K) | 12.19 [42] | ||||
| (ps) | 0.1 [42] | ||||
| (QO)(1019 cm-3) | 3.68 [28] | ||||
| (QO)(1019 cm-3) | 3.42 [28] | - | |||
IV Discussion
Generally, the large MR in WTe2 is discussed in the context of the two-band model [4, 30, 41, 34, 59, 42]. This model, however, suffers from an oversimplification because it clubs together holes from two different pockets with possibly different mobilities into a single band and similarly for electrons from different pockets into another single band. The MR in this model is given by [30, 41, 42]
where is the magnetic field and , , and are total electron concentration, total hole concentration, electron mobility and hole mobility, respectively (see Eqns. S1(a) and S1(b) in the Supplemental Material [50]).
The success of the two-band model lies in the fact that it predicts a parabolic and non-saturating MR () in case of perfect compensation (). In this model, when the charge compensation is imperfect, the MR saturates at high fields, the value of which depends on the mobility. In fact, a very large mobility is counterproductive to a large MR, as it will lead to saturation at low field. For the optimum value of the mobility, the field at which the MR saturates can be beyond that accessible experimentally. However, the two-band model suffers from over parametrization and can lead to wrong conclusions as has been seen in the case of ZrSiSe where a Lifshitz transition inferred from the two-band fit has been later proved to be incorrect [60].
Considering this mutual dependency in the two-band model and the fact that it discards much information about the Fermi surface, it is clear that the measured changes in quantum oscillation frequencies give a more accurate picture of the evolution of the Fermi surface. A simple analysis of quantum oscillations suggests that the charge compensation improves with Mo substitution, at least up to . This last statement relies on the reliability of two assumptions underlying the estimate of the carrier densities: (i) that the prior band-structure calculations are correct in their assignment of the holes and electrons to the different Fermi-surface pockets, and (ii) that the pockets evolve isotropically in all cases, thus allowing the use of the spherical approximation. In any case, the decrease in MR that accompanies the increasing Mo substitution should be attributed solely or in part due to a decrease in transport mobility, as evidenced by the RRR data. It is possible that doping with Mo, while nominally isoelectronic, might affect the stability of other charged defects in the material, and could lead to changes in the Fermi energy and hence the charge compensation in addition to those arising from changes in the lattice parameters. Any such changes will be captured by the evolution of the quantum oscillation frequencies.
While the transport scattering decreases, the quantum-oscillation analysis show that the quantum scattering time remains largely unchanged across the doping range investigated. A similar observation has been reported in Ni substituted WTe2 where increases while decreases with Ni substitution [61]. These effects likely arise from an increase in large-angle scattering due to the introduction of the impurity ions. This will disproportionally affect the transport scattering time, while the quantum scattering time must instead be dominated by small-angle scattering.
In conclusion, we analyse the evolution of the quantum oscillations in WTe2 with Mo substitution. The MR decreases with substitution, but the charge compensation improves based on the quantum oscillation data. We attribute the MR decrease to the decrease in transport mobility, which is evident from the RRR values. Thus, we show that it is possible that the MR in WTe2 can decrease even with a very good level of charge compensation provided there is a fall in the transport mobility. Our quantum oscillation analysis further indicates that the quantum scattering time is unaffected by the Mo substitution. Ref 42 reports changes in mobility in WTe2 nanoflakes on Mo doping, similar to our results. However, they also use the two-band model to conclude that the charge balance remains undisturbed. Our data on bulk crystals shows that this approach is not valid for this material.
V Acknowledgments
We acknowledge the EPSRC, UK for funding this work (Grant Nos. EP/L014963/1 and EP/M028771/1). We thank T. Orton and P. Ruddy for technical assistance. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) licence to any Author Accepted Manuscript version arising from this submission. Data presented in this paper will be made available at XXXXXXX.